Determine the mass in grams of 6.65 × 10²¹ atoms of barium. (The mass of one mole of barium is 137.33 g.)


1
Expert's answer
2021-10-12T02:07:45-0400

Proportion:

A mole of barium contains 6.02×1023 atoms.

Proportion:

1 mol - 6.02×1023 atoms

x mol - 6.65 ×1021 atoms

"x= \frac{1 \times 6.65 \times 10^{21}}{6.02 \times 10^{23}} = 1.104 \times 10^{-2} \;mol \\

m = n \times M \\

m(barium) = 1.104 \times 10^{-2} \times 137.33 = 1.517 \;g"

Answer: 1.517 g